Status Report

NOAA GOES SATOPS Morning Report 11 Jan 2001

By SpaceRef Editor
January 11, 2001
Filed under ,

SATOPS MORNING REPORT – January 11, 2001

POES

NOAA-15 rev 13837W at 1324Z resulted in no data from G4B being recorded
following an IMUX time out and crash.  G4B was recovered 8+ hours
late on rev 13842F at 2147Z.

No change in the status of the POES spacecraft.

 

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.

 

GOES

No change in the status of the GOES spacecraft.

GOES East:  Nominal operations over the past 24 hours.

GOES West:  Rapid scan operations 1733Z-0333Z supporting NWS Monterey
request, Southern California flash flood threat.

Other

Wallops CDA maintenance of switchgear equipment successful.  More
work today.

 

Questions or comments to [email protected]

SpaceRef staff editor.